Let E and B denote electric and amgnetic fields in a frame S and E1 and B1 in another frame S1 moving with respect to S at a velocity v. Which of the folowing equations are wrong?
 
(A) B1y = By + vEz/c2
(B) E1y = Ey  - vBz/c2
(C) B1y = By + vEz
(D) E1y = Ey + vBz
 
please explain the answer. This question is from HCV page 230.


Grade:11

1 Answers

ROSHAN MUJEEB
3 years ago
From Lorentz transformation,
Ey′​=γ(Ey​−vBz​)
By′​=γ(By​+c2v​Ez​)
whereγ=1−c2v2​​1​where the symbols have their usual meaning.
whencv​<<1γ=1, then the the transformation equations will hold.